5. A triangular playground has its sides measured. They sides are f metres, f + 4 metres and f + 5 metres.
a) Write an expression for the perimeter of the playground.
b) The perimeter is 51 metres. Form a suitable equation using this information.
c) What are the lengths of all three sides of the playground

Answer by Theo(13342)

p = perimeter.
p = f + f + 4 + f + 5
combine like terms to get p = 3f + 9
when p = 51, the formula becomes 51 = 3f + 9
subtract 9 from both sides of the equation to get 42 = 3f
divide both sides of the equation by 3 to get 14 = f.
the length of the 3 sides are 14, 14 + 4 = 18, 14 + 5 = 19
their sum is equal to 14 + 18 + 19 = 51.